# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 2. Preliminaries
In the case where $g=3$, we will show that, for any positive real number $\lambda$, there is an integer $n$ such that the dilatation of $\rho_{n}$ is more than $\lambda$. Let $\tau$ be the train track on $\Sigma_{3}$ which carries each element of $\mathscr{C}\cup\mathscr{D}$. As shown Figure 4, the train track $\tau$ has 27 branches $\delta_{i}$, for $i=1,\ldots,9$, $\mu_{3j-2}^{k}$, for $j=1,2,3,k=1,2,3$, $\mu_{3j-1}$, for $j=1,2,3$ and $\mu_{3j}^{k}$, for $j=1,2,3,k=1,2$, where the transverse measure for each branch is indicated by the same symbol. The train track $\rho_{n}^{3}(\tau)$ is carried by $\tau$, and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}(\tau)$ are also carried by $\tau$ for $i=1,\ldots,g$. The set $H$ of transverse mesures on $\tau$ determined by the equations $\mu_{3j-2}^{1}=\mu_{3j-2}^{2}=\mu_{3j-2}^{3},\mu_{3j}^{1}=\mu_{3j}^{2}$ is preserved by the action of $t_{a_{i}}$ $t_{b_{i}}^{-1}$ and $t_{c_{i}}$. We denote $\mu_{3j-2}^{1},\mu_{3j-2}^{2},\mu_{3j-2}^{3}$ by $\mu_{3j-2}$ and $\mu_{3j}^{1},\mu_{3j}^{2}$ by $\mu_{3j}$ for short. The actions of $\rho_{n}^{3}$ and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}$ for $i=1,\ldots,3$ on $H$ are determined by the actions of them on values $\mu_{j}$, for $j=1,\ldots,9$. For the homeomorphism $\varphi$ on $\Sigma_{3}$ such that $\varphi(\tau)$ is carried by $\tau$, let $M_{\varphi}$ be the incident matrix for $\varphi$, that is, the $(i,j)$ entry $(M_{\varphi})_{i,j}$ of $M_{\varphi}$ is the times $\varphi(\eta_{j})$ passing through $\mu_{i}$. Then the matrix $M_{\rho_{n}^{3}}$ is a non-negative irreducible matrix and the maximal eigenvalue $\mathrm{PF}(M_{\rho_{n}^{3}})$ of $M_{\rho_{n}^{3}}$ equals the dilatation $\lambda_{n}^{3}$ of $\rho_{n}^{3}$, where $\lambda_{n}$ is the dilatation of $\rho_{n}$. Let $M$ and $N$ be matrices with the same size. If, for each $(i,j)$ entries $M_{i,j}$ and $N_{i,j}$ of $M$ and $N$ respectively, $M_{i.j}\geq N_{i,j}$, then we write $M\geq N$. If $\psi=t_{a_{i}}$ or $t_{b_{i}}^{-1}$ or $t_{c_{i}}$, then $M_{\psi}=I_{9}+P$, where $I_{9}$ is the identity matrix of order 9 and $P$ is a matrix all of whose entries are non-negative. Therefore,
$M_{\rho_{n}^{3}}\geq M_{t_{b_{1}}^{-n}}M_{t_{c_{1}}}$ $=\left(\begin{array}[]{ccc|ccc|ccc}1&n&n&0&0&0&0&0&n\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)\left(\begin{array}[]{ccc|ccc|ccc}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 1&0&1&1&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)$
$=\left(\begin{array}[]{ccc|ccc|ccc}n+1&n&n&n&0&0&0&0&n\\ 0&1&0&0&0&0&0&0&0\\ 1&0&1&1&0&0&0&0&0\\ \hline\cr 0&0&1&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right)\geq\left(\begin{array}[]{ccc|ccc|ccc}n+1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ \hline\cr 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ \hline\cr 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{array}\right).$
@@
It is well-known that, if $M$ is a non-negative irreducible matrix and $N$ is a non-negative matrix such that $M\geq N$, then every eigenvalue $\sigma$ of $N$ satisfies $\mathrm{PF}(M)\geq|\sigma|$ (see Theorem 9.1.1 of [24], for example). Therefore, $\lambda_{n}^{3}\geq n+1$.
We can show the same result for the case where $g\geq 4$, as above. ∎
@@
The element $\mathcal{R}^{-1}$ is called a root of Dehn twist of degree $12g-$ 1 (see [22, 21, 18]).
@@
Note that $\mathcal{R}$ is in $\mathcal{H}_{g}$ since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$. It is well-know that the relation $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{2g-1}=t_{\alpha_{2g+1}}^{2}$ holds in $\mathcal{H}_{g}$ (see Section 4.4.1 in [6]). This relation is called the chain relation. Since $\alpha_{2g+1}$ is disjoint from $\alpha_{1},\ldots,\alpha_{2g-1}$, by the chain relation, we have
@@
In this section, we prove Theorems 1.1 and  1.4 in the introduction.
Recall that the notations $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1}$ and $\beta$ denote the simple closed curves on $\Sigma_{g}$ as in Figure 3. Note that $\alpha_{i}$ (resp. $\beta$) intersects $\alpha_{i+1}$ (resp. $\alpha_{4}$) transversely at exactly one point for $i=1,2,\ldots,2g-1$. Moreover, $\alpha_{i}$ (resp. $\beta$) is disjoint from $\alpha_{j}$ if $|i-j|>1$ (resp. $j\neq 4$). In this section, we will use the following relations
$t_{t_{\alpha_{i+1}}^{-1}(\alpha_{i})}=t_{t_{\alpha_{i}}(\alpha_{i+1})}$ $\mathrm{and}$ $t_{t_{\alpha_{i+1}}(\alpha_{i})}=t_{t_{\alpha_{i}}^{-1}(\alpha_{i+1})}.$
@@
Note that if a Dehn twist $t_{c}$ about a simple closed curve $c$ and an element $f$ are in $H$, then $t_{f}(c)$ is also in $H$ by the relation $t_{f(c)}=ft_{c}f^{-1}$.
We have $t_{\alpha_{1}}=h_{1}h_{2}^{-1}\in H$, and hence $t_{h_{2}(\alpha_{1})}$ is in $H$ by $h_{2}\in H$. Since $h_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})$ and $t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we obtain
$t_{h_{2}(\alpha_{1})}=t_{t_{\alpha_{2}}^{-1}(\alpha_{1})}=t_{t_{\alpha_{1}}(\alpha_{2})}=t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{1}}^{-1},$
and therefore, $t_{\alpha_{1}}t_{\alpha_{2}}t_{\alpha_{1}}^{-1}$ is in $H$. By $t_{\alpha_{1}}\in H$, we have $t_{\alpha_{2}}\in H$. We set $h_{3}=t_{\alpha_{2}}h_{2}$, then $h_{3}$ is in $H$, and hence $t_{h_{3}(\alpha_{2})}$ is in $H$. Since $h_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})$ and $t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we get
$H\ni t_{h_{3}(\alpha_{2})}=t_{t_{\alpha_{3}}(\alpha_{2})}=t_{t_{\alpha_{2}}^{-1}(\alpha_{3})}=t_{\alpha_{2}}^{-1}t_{\alpha_{3}}t_{\alpha_{2}},$
and therefore $t_{\alpha_{2}}^{-1}t_{\alpha_{3}}t_{\alpha_{2}}$ is in $H$. By $t_{\alpha_{2}}\in H$, we have $t_{\alpha_{3}}\in H$. Repeating this argument, we can show that $t_{\alpha_{1}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H$. Since $\mathcal{H}_{g}$ is generated by these Dehn twists, we see that $H=\mathcal{H}_{g}$, and the proof is complete. ∎
@@
Lemmas 3.3 and 3.4 below are used to show Theorem 3.2. In the proofs of Lemmas 3.3 and 3.4 and Theorem 3.2, we will repeatedly use the fact that if a Dehn twist $t_{c}$ about a simple closed curve $c$ and an element $f$ are in $G$, then $t_{f}(c)$ is also in $G$ by the relation $t_{f(c)}=ft_{c}f^{-1}$.
@@
For $g\geq 3$, the elements $t_{\alpha_{1}},t_{\alpha_{2}},t_{\alpha_{3}},t_{t_{\alpha_{4}}(\alpha_{5})},t_{t_{\alpha_{5}}(\alpha_{6})},f_{3},f_{4}$ and $f_{5}$ are contained in $G$.
@@
Note that we have $t_{\alpha_{1}}=t_{\beta}t_{\alpha_{1}}t_{\beta}^{-1}=f_{1}f_{2}^{-1}$ (since $\alpha_{1}$ is disjoint from $\beta$). This means that $t_{\alpha_{1}}$ is in $G$, and hance $t_{\alpha_{1}}^{-1}f_{2}$ is also in $G$. Since we find that $f_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we have $t_{\alpha_{1}}^{-1}f_{2}(\alpha_{1})=\alpha_{2}$. Therefore, we get
$t_{\alpha_{2}}\in G$
by $t_{\alpha_{1}},t_{\alpha_{1}}^{-1}f_{2}\in G$. Moreover, by $f_{3}=t_{\alpha_{2}}f_{2}$ (from the fact that $\alpha_{2}$ is disjoint from $\beta$), we get $f_{3}\in G$. Since we find that $f_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we have $t_{\alpha_{2}}f_{3}(\alpha_{2})=\alpha_{3}$, and therefore we see that
$t_{\alpha_{3}}\in G$
by $t_{\alpha_{2}},t_{\alpha_{2}}f_{3}\in G$. From this, we get $f_{4}=t_{\alpha_{3}}^{-1}f_{3}\in G$ (since $\alpha_{3}$ is disjont from $\beta$). Here, since $\alpha_{3}$ (resp. $\alpha_{4}$) intersects $\alpha_{4}$ (resp. $\beta$) transversely at exactly one point, we find that $f_{4}(\alpha_{3})=t_{\beta}t_{\alpha_{4}}^{-1}(\alpha_{3})=t_{\beta}t_{\alpha_{3}}(\alpha_{4})=t_{\alpha_{3}}t_{\beta}(\alpha_{4})$ (since $\alpha_{3}$ is disjoint from $\beta$). This gives $t_{\alpha_{3}}^{-1}f_{4}(\alpha_{3})=t_{\beta}(\alpha_{4})$, and hence we have
$t_{t_{\beta}(\alpha_{4})}\in G$
by $t_{\alpha_{3}},t_{\alpha_{3}}^{-1}f_{4}\in G$. Furthermore, the following holds:
$G\ni f_{4}$ $=t_{\beta}t_{\alpha_{4}}^{-1}\cdot t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{t_{\beta}(\alpha_{4})}f_{5}.$
This means that $f_{5}\in G$. Since $\beta$ and $\alpha_{4}$ are disjont from $\alpha_{6},\ldots,\alpha_{2g}$, we have $f_{5}(t_{\beta}(\alpha_{4}))=t_{\beta}t_{\alpha_{5}}(t_{\beta}(\alpha_{4}))$. Note that $\beta$ is also disjoint from $\alpha_{5}$. Therefore, using the braid relation $t_{\alpha_{5}}t_{\beta}=t_{\beta}t_{\alpha_{5}}$, the relations $t_{\alpha_{5}}(\alpha_{4})=t_{\alpha_{4}}^{-1}(\alpha_{5})$, $t_{\beta}(\alpha_{5})=\alpha_{5}$ and the braid relation $t_{\beta}t_{\alpha_{4}}t_{\beta}=t_{\alpha_{4}}t_{\beta}t_{\alpha_{4}}$, we have
$t_{t_{\beta}(\alpha_{4})}f_{5}(t_{\beta}(\alpha_{4}))$ $=t_{t_{\beta}(\alpha_{4})}t_{\beta}^{2}t_{\alpha_{5}}(\alpha_{4})$
$=t_{t_{\beta}(\alpha_{4})}t_{\beta}^{2}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{4}}t_{\beta}^{-1}t_{\beta}^{2}t_{\alpha_{4}}^{-1}(\alpha_{5})$
$=t_{\beta}t_{\alpha_{4}}t_{\beta}t_{\alpha_{4}}^{-1}(\alpha_{5})$
@@
$=t_{\alpha_{4}}(\alpha_{5}).$
This gives $t_{t_{\alpha_{4}}(\alpha_{5})}=t_{t_{t_{\beta}(\alpha_{4})}f_{5}(t_{\beta}(\alpha_{4}))}(=(t_{t_{\beta}(\alpha_{4})}f_{5})t_{t_{\beta}(\alpha_{4})}(t_{t_{\beta}(\alpha_{4})}f_{5})^{-1})$, and hence we have
$t_{t_{\alpha_{4}}(\alpha_{5})}\in G$
by $t_{t_{\beta}(\alpha_{4})},t_{t_{\beta}(\alpha_{4})}f_{5}\in G$. Since $\alpha_{i}$ is disjoint from $\alpha_{j}$ if $|i-j|\geq 2$, using the relation $t_{\alpha_{6}}^{-1}(\alpha_{5})=t_{\alpha_{5}}(\alpha_{6})$ and the braid relation $t_{\alpha_{5}}t_{\alpha_{4}}t_{\alpha_{5}}=t_{\alpha_{4}}t_{\alpha_{5}}t_{\alpha_{4}}$, we have
$f_{3}(t_{\alpha_{4}}(\alpha_{5}))$ $=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}(t_{\alpha_{4}}(\alpha_{5}))$
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{4}}(\alpha_{5})$
@@
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{5}}(\alpha_{6}).$
Since $\beta$ and $\alpha_{3}$ are disjoint from $\alpha_{5}$ and $\alpha_{6}$, we obtain $f_{3}(t_{\alpha_{4}}(\alpha_{5}))=t_{\alpha_{5}}(\alpha_{6})$, and hence we have
$t_{t_{\alpha_{5}}(\alpha_{6})}\in G.$
by $t_{t_{\alpha_{4}}(\alpha_{5})},f_{3}\in G$, and the proof is complete. ∎
@@
Let $g\geq 4$, and set $F_{2k-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2k-3}}f_{2k-1}$ for $3\leq k\leq g$. We assume that the elements $t_{t_{\alpha_{2i-1}}(\alpha_{2i})},F_{2i-1},F_{2i+1}$ are contained in $G$ for some $i$ such that $13\leq i\leq g-$ 1, then we have
@@
    $t_{t_{\alpha_{2i}}(\alpha_{2i+1})},t_{t_{\alpha_{2i+1}}(\alpha_{2i+2})}\in G$ if $13\leq i\leq g-$ 1,

### v3

@@ 2. Preliminaries
We will show that, for any positive real number $\lambda$, there is an integer $n$ such that the dilatation of $\rho_{n}$ is more than $\lambda$. Let $\tau$ be the train track on $\Sigma_{g}$ which carries each element of $\mathscr{C}\cup\mathscr{D}$. As shown in Figure 4, the train track $\tau$ has branches $\delta_{i}$ for $i=1,\ldots,3g$, $\mu_{3j-2}^{k}$ for $j=1,\ldots,g,k=1,2,3$, $\mu_{3j-1}$ for $j=1,\ldots,g$ and $\mu_{3j}^{k}$ for $j=1,\ldots,g,k=1,2$. For each branch, the transverse measure is indicated by the same symbol. The train track $\rho_{n}^{g}(\tau)$ is carried by $\tau$, and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}(\tau)$ are also carried by $\tau$ for $i=1,\ldots,g$. The set $H$ of transverse mesures on $\tau$ determined by the equations $\mu_{3j-2}^{1}=\mu_{3j-2}^{2}=\mu_{3j-2}^{3},\mu_{3j}^{1}=\mu_{3j}^{2}$ is preserved by the action of $t_{a_{i}}$ $t_{b_{i}}^{-1}$ and $t_{c_{i}}$. We denote $\mu_{3j-2}^{1},\mu_{3j-2}^{2},\mu_{3j-2}^{3}$ by $\mu_{3j-2}$ and $\mu_{3j}^{1},\mu_{3j}^{2}$ by $\mu_{3j}$ for short. The actions of $\rho_{n}^{g}$ and $t_{a_{i}}t_{b_{i}}^{-n}t_{c_{i}}$ for $i=1,\ldots,g$ on $H$ are determined by the actions of them on values $\mu_{j}$ for $j=1,\ldots,3g$. For the homeomorphism $\varphi$ on $\Sigma_{g}$ such that $\varphi(\tau)$ is carried by $\tau$, let $M_{\varphi}$ be the incident matrix for $\varphi$, that is, the $(i,j)$ entry $(M_{\varphi})_{i,j}$ of $M_{\varphi}$ is the times $\varphi(\mu_{j})$ passing through $\mu_{i}$. Then the matrix $M_{\rho_{n}^{g}}$ is a non-negative irreducible matrix and the maximal eigenvalue $\mathrm{PF}(M_{\rho_{n}^{g}})$ of $M_{\rho_{n}^{g}}$ equals the dilatation $\lambda_{n}^{g}$ of $\rho_{n}^{g}$, where $\lambda_{n}$ is the dilatation of $\rho_{n}$. Let $M$ and $N$ be matrices with the same size. If, for each $(i,j)$ entries $M_{i,j}$ and $N_{i,j}$ of $M$ and $N$ respectively, $M_{i.j}\geq N_{i,j}$, then we write $M\geq N$. If $\psi=t_{a_{i}}$ or $t_{b_{i}}^{-1}$ or $t_{c_{i}}$, then $M_{\psi}=I_{3g}+P$, where $I_{3g}$ is the identity matrix of order $3g$ and $P$ is a matrix all of whose entries are non-negative. Therefore,
$M_{\rho_{n}^{g}}\geq M_{t_{b_{1}}^{-n}}M_{t_{c_{1}}}$
$=\left(\begin{array}[]{c|c|c|c}{\begin{array}[]{ccc}1&n&n\\ 0&1&0\\ 0&0&1\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}0&0&n\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr O&O&{\begin{array}[]{c}\\ I_{3g-9}\\ \\ \end{array}}&O\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}\end{array}\right)\left(\begin{array}[]{c|c|c|c}{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 1&0&1\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 1&0&0\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr O&O&{\begin{array}[]{c}\\ I_{3g-9}\\ \\ \end{array}}&O\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}\end{array}\right)$
$=\left(\begin{array}[]{c|c|c|c}{\begin{array}[]{ccc}n+1&n&n\\ 0&1&0\\ 1&0&1\end{array}}&{\begin{array}[]{ccc}n&0&0\\ 0&0&0\\ 1&0&0\end{array}}&O&{\begin{array}[]{ccc}0&0&n\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr O&O&{\begin{array}[]{c}\\ I_{3g-9}\\ \\ \end{array}}&O\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}\end{array}\right)\geq\left(\begin{array}[]{c|c|c|c}{\begin{array}[]{ccc}n+1&0&0\\ 0&1&0\\ 0&0&1\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}&O&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}\\ \hline\cr O&O&{\begin{array}[]{c}\\ I_{3g-9}\\ \\ \end{array}}&O\\ \hline\cr{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&{\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\end{array}}&O&{\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}}\end{array}\right).$
@@
It is well-known that, if $M$ is a non-negative irreducible matrix and $N$ is a non-negative matrix such that $M\geq N$, then every eigenvalue $\sigma$ of $N$ satisfies $\mathrm{PF}(M)\geq|\sigma|$ (see Theorem 9.1.1 of [24], for example). Therefore, $\lambda_{n}^{g}\geq n+1$. ∎
@@
The element $\mathcal{R}^{-1}$ is called a root of Dehn twist of degree 2g12g-1 (see [22, 21, 18]).
@@
Note that $\mathcal{R}$ is in $\mathcal{H}_{g}$ since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$. It is well-known that the relation $(t_{\alpha_{1}}^{2}t_{\alpha_{2}}t_{\alpha_{3}}\cdots t_{\alpha_{2g-1}})^{2g-1}=t_{\alpha_{2g+1}}^{2}$ holds in $\mathcal{H}_{g}$ (see Section 4.4.1 in [6]). This relation is called the chain relation. Since $\alpha_{2g+1}$ is disjoint from $\alpha_{1},\ldots,\alpha_{2g-1}$, by the chain relation, we have
@@
In this section, we prove Theorems 1.1 and  1.4 in the introduction. We will repeatedly use the fact that if a Dehn twist $t_{c}$ about a simple closed curve $c$ and an element $f$ are in a certain subgroup $G$ of $\mathcal{M}_{g}$, then $t_{f}(c)$ is also in $G$ by the relation $t_{f(c)}=ft_{c}f^{-1}$. Recall that the notations $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1}$ and $\beta$ denote the simple closed curves on $\Sigma_{g}$ as in Figure 3. Note that $\alpha_{i}$ (resp. $\beta$) intersects $\alpha_{i+1}$ (resp. $\alpha_{4}$) transversely at exactly one point for $i=1,2,\ldots,2g-1$. Moreover, $\alpha_{i}$ (resp. $\beta$) is disjoint from $\alpha_{j}$ if |ij|>1|i-j|>1 (resp. $j\neq 4$). In this section, we will use the following fact:
$t_{\alpha_{i+1}}^{-1}(\alpha_{i})=t_{\alpha_{i}}(\alpha_{i+1}),$ $\mathrm{and}$ $t_{\alpha_{i+1}}(\alpha_{i})=t_{\alpha_{i}}^{-1}(\alpha_{i+1}).$
We denote by $\gamma_{i}$ the simple closed curve $t_{\alpha_{i}}(\alpha_{i+1})$ for $i=4,5,\ldots,2g-1$.
@@
We see that $t_{\alpha_{1}}=h_{1}h_{2}^{-1}$ is in $H$, and hence $t_{h_{2}(\alpha_{1})}$ is in $H$. Since $h_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we obtain $t_{h_{2}(\alpha_{1})}=t_{t_{\alpha_{1}}(\alpha_{2})}$ and therefore, $t_{\alpha_{2}}$ is in $H$. We set $h_{3}=t_{\alpha_{2}}h_{2}$, then $h_{3}$ is in $H$, and hence $t_{h_{3}(\alpha_{2})}$ is in $H$. By $h_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we have $t_{h_{3}(\alpha_{2})}=t_{t_{\alpha_{2}}^{-1}(\alpha_{3})}$, and therefore $t_{\alpha_{3}}$ is in $H$. Repeating this argument, we can show that $t_{\alpha_{1}},\ldots,t_{\alpha_{2g+1}}$ are contained in $H$. Since $\mathcal{H}_{g}$ is generated by these Dehn twists, we see that $H=\mathcal{H}_{g}$, and the proof is complete. ∎
@@
Lemmas 3.3 and 3.4 below are used to show Theorem 3.2.
@@
For $g\geq 3$, the elements $t_{\alpha_{1}},t_{\alpha_{2}},t_{\alpha_{3}},t_{\gamma_{4}},t_{\gamma_{5}},f_{3},f_{4}$ and $f_{5}$ are contained in $G$.
@@
By $t_{\alpha_{1}}=t_{\beta}t_{\alpha_{1}}t_{\beta}^{-1}=f_{1}f_{2}^{-1}$, $t_{\alpha_{1}}$ is in $G$, and hence $t_{\alpha_{1}}^{-1}f_{2}$ is also in $G$. Since we find that $f_{2}(\alpha_{1})=t_{\alpha_{2}}^{-1}(\alpha_{1})=t_{\alpha_{1}}(\alpha_{2})$, we have $t_{\alpha_{1}}^{-1}f_{2}(\alpha_{1})=\alpha_{2}$. Therefore, $t_{\alpha_{2}}$ is in $G$. By $f_{3}=t_{\alpha_{2}}f_{2}$, we see that $f_{3}$ is in $G$. Since we find that $f_{3}(\alpha_{2})=t_{\alpha_{3}}(\alpha_{2})=t_{\alpha_{2}}^{-1}(\alpha_{3})$, we have $t_{\alpha_{2}}f_{3}(\alpha_{2})=\alpha_{3}$, and therefore we see that $t_{\alpha_{3}}$ is in $G$. From this, $f_{4}=t_{\alpha_{3}}^{-1}f_{3}$ is in $G$. Note that we find that $f_{4}(\alpha_{3})=t_{\beta}t_{\alpha_{4}}^{-1}(\alpha_{3})=t_{\beta}t_{\alpha_{3}}(\alpha_{4})=t_{\alpha_{3}}t_{\beta}(\alpha_{4})$. This gives $t_{\alpha_{3}}^{-1}f_{4}(\alpha_{3})=t_{\beta}(\alpha_{4})$, and hence $t_{\delta_{4}}$ is in $G$, where $\delta_{4}$ is the simple closed curve $t_{\beta}(\alpha_{4})$. Furthermore, the following holds:
$f_{4}$ $=t_{\beta}t_{\alpha_{4}}^{-1}\cdot t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{7}}t_{\alpha_{8}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
$=t_{\delta_{4}}^{-1}f_{5}.$
This means that $f_{5}$ is in $G$. Since $\beta$ and $\alpha_{4}$ are disjont from $\alpha_{6},\ldots,\alpha_{2g}$, we obtain $f_{5}(\delta_{4})=t_{\beta}t_{\alpha_{5}}(\delta_{4})$. This gives
$t_{\delta_{4}}f_{5}(\delta_{4})$ $=t_{\beta}t_{\alpha_{4}}t_{\beta}^{-1}t_{\beta}t_{\alpha_{5}}t_{\beta}(\alpha_{4})$
$=t_{\beta}t_{\alpha_{4}}t_{\beta}t_{\alpha_{5}}(\alpha_{4})$
@@
$=t_{\alpha_{4}}(\alpha_{5})(=\gamma_{4}).$
Therefore, $t_{\gamma_{4}}$ is in $G$. Moreover, we have
$f_{3}(\gamma_{4})$ $=t_{\beta}t_{\alpha_{3}}t_{\alpha_{4}}^{-1}t_{\alpha_{5}}t_{\alpha_{6}}^{-1}t_{\alpha_{4}}(\alpha_{5})$
@@
$=t_{\beta}t_{\alpha_{3}}t_{\alpha_{5}}(\alpha_{6})$
$=t_{\alpha_{5}}(\alpha_{6})(=\gamma_{5}).$
Therefore, $t_{\gamma_{5}}$ is in $G$, and the proof is complete. ∎
@@
Let $g\geq 4$, and set $F_{2k-1}=t_{\alpha_{3}}t_{\alpha_{5}}\cdots t_{\alpha_{2k-3}}f_{2k-1}$ for $3\leq k\leq g$. We assume that the elements $t_{\gamma_{2i-1}},F_{2i-1},F_{2i+1}$ are contained in $G$ for some $i$ such that $3\leq i\leq g-1$, then the following holds:
@@
    $t_{\gamma_{2i}}$ and $t_{\gamma_{2i+1}}$ are in $G$ if $3\leq i\leq g-1$,